# Pinned distances of planar sets with low dimension

### v1

@@ 1. Introduction
A fundamental problem in geometric measure theory is to relate the (Hausdorff) dimension of $E$ and the dimension of its distance set $\Delta(E)$. Falconer conjectured [7] that, if $E$ is analytic and $\dim_{H}(E)>n/2$, then $\Delta(E)$ has positive Lebesgue measure (and so $\dim_{H}(\Delta(E))=1$). This is still unproven in all dimensions $n\geq 2$.
A generalization of this is the pinned distance problem. If $E\subseteq\mathbb{R}^{n}$ and $x\in\mathbb{R}^{n}$, the pinned distance set of $E$ with respect to $x$ is the set
@@
The dimension version of Falconer’s conjecture for pinned distance sets is the following.
@@
Let $E\subseteq\mathbb{R}^{n}$ be an analytic set with $\dim_{H}(E)>n/2$. Then
@@
Substantial progress has been made on this question in the past few years [6, 9, 3, 4, 5].
A full resolution of Conjecture 1 problem seems to be beyond current techniques, even in the $n=2$ case. However, recent work has proven strong lower bounds on the dimension of pinned distance sets in the plane [23, 24, 11, 27]. If $E\subseteq\mathbb{R}^{2}$ is analytic, and $d:=\dim_{H}(E)>1$, the current records are
@@
Unfortunately, the techniques in both [12] and [8] rely crucially on the assumption that $\dim_{H}(E)>1$.
In this paper, we prove lower bounds on the pinned distance sets of planar sets of small dimension. There is significantly less known in this setting. A fundamental issue had been the lack of progress on radial projections for low-dimensional sets. Nevertheless, Shmerkin and Wang gave the first strong progress on this problem [^1] A key ingredient of Shmerkin and Wang’s proof was a result on radial projections of low dimensional sets. This component was then improved in [22].. They proved [25] that, for any analytic set $E\subseteq\mathbb{R}^{2}$ with $d:=\dim_{H}(E)\leq 1$,
Moreover, they proved that if $E$ is weakly regular, i.e., $\dim_{H}(E)=\dim_{P}(E)$, then $\sup\limits_{x\in E}\dim_{H}(\Delta_{x}E)=d$. In [5], Du, Ou, Ren and Zhang proved that
(2) $\sup\limits_{x\in E}\dim_{H}(\Delta_{x}E)\geq\frac{5}{3}d-1,$
which improves on (1) when $d$ is sufficiently close to one.
In this paper, we establish stronger bounds on the dimension of pinned distance sets, for certain ranges of $d$. Our main theorem is the following.
@@
Note that the bound improves as $E$ becomes more regular, i.e., as the packing dimension of $E$ approaches its Hausdorff dimension. An important special case (when $D=2$) is captured by the following corollary.
@@
Suppose $E\subseteq\mathbb{R}^{2}$ is analytic set and $d:=\dim_{H}(E)\leq 1$. Then,
@@
The proof of this theorem uses effective methods, and follows the general framework of the authors previous paper [8]. The assumption that $\dim_{H}(E)<1$ necessitates important changes to previous work.
One of the most crucial differences is the use of the recent breakthrough by Orponen, Shmerkin and Wang [22] on radial projections for low-dimensional sets. In the language of effective dimension, this allows us to find $x,y\in E$ whose direction $e:=\frac{y-x}{|x-y|}$ is sufficiently complex. Additionally, a number of the algorithmic tools of [8] need to be adapted to this setting.
@@
We can strengthen this notion by requiring the existence of specific points in $X$ achieving the desired bound. Hence, we say that $X$ is universal for pinned distances for $\mathcal{C}$ if, for every $Y\in\mathcal{C}$ there exists some $x\in X$ such that
@@
Unpacking the definitions, this theorem shows that there is a fixed set $X\subseteq\mathbb{R}^{2}$ such that, for every $Y$ with $\dim_{H}(Y)<\dim_{H}(X)$, $\sup\limits_{x\in X}\dim_{H}(\Delta_{x}(Y))$ is maximal. Moreover, combined with a previous theorem of the authors [8], we have the following corollary.
@@
In particular, if we take any self-similar four-corner Cantor set $C$ of dimension strictly greater than $1$, then for every Borel set $Y\subseteq\mathbb{R}^{2}$, there exists some $x\in C$ such that
@@ 3.3. Main projection theorem
Case $S=1$: Suppose the green portion of our red-green-blue sequence starts at precision $r_{1}$ and ends at precision $r_{1}+s$, for some $s\geq t$. Let $r_{2}\geq r_{1}+s$ be the minimal precision such that $[r_{2},r]$ is the union of yellow intervals [^2] No such precision necessarily exists. In this case, set $r_{2}=r$..
@@ 5. Reduction to the classical bounds
We note that $e^{\bot}$ and $e$ are Lipschitz-bi-computable, and so $K_{r}(e)=K_{r}(e^{\bot})+O(\log r)$. By a pointwise analog of Marstrand’s projection theorem, [^3] Essentially proven in [17], but also a result of the projection bound in section 3 arising from a one interval partition
@@ 6. Universal pin sets
In the second subsection, we assume additional regularity in $X$, namely that it is AD regular [^4] Full Ahlfors-David regularity isn’t necessary, it is possible to weaken the assumption slightly. However, we will require AD regularity to simplify the exposition.. In this case, we are guaranteed the existence of points that are extremely regular, with their complexity functions only deviating from $\dim_{H}(x)r$ by a logarithmic term. This enables us to show an optimal effective theorem. We only consider the case that $\dim_{H}(X)>1$ so we can employ a somewhat different reduction, though one that still depends on [22].

### v2

@@ 1. Introduction
A fundamental problem is to relate the sizes of $E$ and $\Delta(E)$; if $E$ is large, it should contain “many” different distances. For finite sets $E$, this is the Erdős distinct distances problem, which was essentially solved in the plane by Guth and Katz [10]. In the continuum setting, Falconer conjectured [7] that, if $E$ is Borel and has Hausdorff dimension more than $n/2$, then $\Delta(E)$ has positive Lebesgue measure. Falconer’s conjecture is still open in all dimensions $n\geq 2$. For $E$ no larger than the $n/2$ threshold, the distance sets do not necessarily have positive measure, but one still expects there to be some lower bound on the Hausdorff dimension of $\Delta(E)$.
Working in the plane, we improve upon the best known lower bounds of this type when .$1.5<\dim_{H}(E)<$ 1. In particular, we strengthen bounds on the size of “pinned” distance sets. If $E\subseteq\mathbb{R}^{n}$ and $x\in\mathbb{R}^{n}$, the pinned distance set of $E$ with respect to $x$ is
@@
Note that if $x\in E$, then $\Delta_{x}(E)\subseteq\Delta(E)$, so appropriate lower bounds on the size of pinned distance sets imply the same lower bounds for distance sets. The dimension version of Falconer’s conjecture for pinned distance sets is the following:
@@
Let $E\subseteq\mathbb{R}^{n}$ be a Borel set with $\dim_{H}(E)>n/2$. Then
@@
Although a full resolution of Conjecture 1 seems to be beyond current techniques, substantial progress has been made on this question in the past few years [6, 9, 3, 4, 5]. Moreover, recent work has proven strong lower bounds on the dimension of pinned distance sets in the plane [25, 26, 12, 29]. If $E\subseteq\mathbb{R}^{2}$ is analytic (i.e. the continuous image of a Borel set), and $d:=\dim_{H}(E)>1$, the current records are
@@
However, the techniques in both [13] and [8] rely crucially on the assumption that $\dim_{H}(E)>1$. In this paper, we establish improved bounds on the dimension of the pinned distance sets in the absence of this assumption. Let $\dim_{P}(E)$ denote the packing dimension of $E$ (another notion of fractal dimension that is dual to Hausdorff dimension and always satisfies $\dim_{P}(E)\geq\dim_{H}(E)$). Our main theorem is the following:
@@
Note that the bound improves as $E$ becomes more regular, i.e., as the packing dimension of $E$ approaches its Hausdorff dimension. Hence, setting $D=2$ gives the following corollary.
@@
Suppose $E\subseteq\mathbb{R}^{2}$ is analytic and $d:=\dim_{H}(E)\leq 1$. Then,
@@
Until recently, less progress had been made on the study of pinned distance sets in this setting than above the $n/2$ threshold. A fundamental issue had been the need for improved bounds on the size of radial projections for small sets. Shmerkin and Wang gave the first strong progress on this problem, proving [27] that, for any analytic set $E\subseteq\mathbb{R}^{2}$ with $d:=\dim_{H}(E)\leq 1$,
Moreover, they proved that if $E$ is weakly regular, i.e., $\dim_{H}(E)=\dim_{P}(E)$, then $\sup\limits_{x\in E}\dim_{H}(\Delta_{x}E)=d$. A key ingredient of Shmerkin and Wang’s proof was a result on radial projections of low-dimensional sets which was then improved in [24]. In [5], Du, Ou, Ren and Zhang proved that
(2) $\sup\limits_{x\in E}\dim_{H}(\Delta_{x}E)\geq\frac{5}{3}d-1,$
which exceeds (1) when $d$ is sufficiently close to one. Corollary 3 improves on both of these bounds provided that $d$ is greater than about .477.477.477.477. Furthermore, the lower bound is $3/4$ when $d=1$, which agrees with the lower bound in [8] and eliminates the discontinuity that had existed at this threshold value.
The proof of our main theorem proceeds via “effective” methods. In particular, we study the algorithmic information content of points as well as the distances and directions between them. Appropriate bounds on the complexity of distances $|x-y|$ imply bounds on the size of the pinned distance set $\Delta_{x}(E)$ via the point-to-set principle of Lutz and Lutz [15]. The proof broadly follows a similar framework to the authors’ previous paper [8]. However, the assumption that $\dim_{H}(E)\leq 1$ necessitates new techniques relative to previous work.
One of the most crucial differences is the use of the recent breakthrough by Orponen, Shmerkin and Wang [24] on radial projections for low-dimensional sets. In the language of effective dimension, this allows us to find $x,y\in E$ whose direction $e:=\frac{y-x}{|x-y|}$ is sufficiently complex. Making use of this theorem requires a new reduction, i.e. a proof that the main theorem follows from its effective analog. Additionally, versions of many of the algorithmic tools of [8] need to be established in this setting, and a careful usage of these tools requires leveraging certain bounds only relevant when the information content of objects is low, as in the case of small sets.
1.1. Universal sets for pinned distances
@@
Note that the right hand side is as large as the pinned distance set can possibly be in the sense of Hausdorff dimension. We can strengthen this notion by requiring the existence of specific points in $X$ achieving the desired bound. Hence, we say that $X$ is universal for pinned distances for $\mathcal{C}$ if, for every $Y\in\mathcal{C}$ there exists some $x\in X$ such that
@@
Unpacking the definitions, this theorem shows that there is a fixed set $X\subseteq\mathbb{R}^{2}$ such that, for every $Y$ with $\dim_{H}(Y)<\dim_{H}(X)$, $\sup\limits_{x\in X}\dim_{H}(\Delta_{x}(Y))$ is maximal. This can be thought of as strengthening the result in [27] that if $Y$ is weakly regular, then it has pinned distance sets of the maximum possible size. We are able to obtain the same conclusion using only regularity in the pin set $X$. Moreover, combined with a previous theorem of the authors (Theorem 7 in [8], which handles the case that $Y$ is large), we have the following corollary.
@@
For instance, if we take any self-similar four-corner Cantor set $C$ of dimension strictly greater than $1$, then for every Borel set $Y\subseteq\mathbb{R}^{2}$, there exists some $x\in C$ such that
@@
Considering again the “unpinned” distance set problem gives another interpretation of Theorem 6. Sets such as $C$ have the property that for any Borel $Y$, one can choose $x\in C$ such that
$\dim_{H}(\Delta(Y\cup\{x\}))\geq\min\{\dim_{H}(Y),1\}$
In particular, (the dimension version of) Falconer’s conjecture is true in the plane for any Borel $Y$, provided that we are allowed to add one point drawn from a small fixed set to $Y$.
@@ 3.3. Main projection theorem
Case $S=1$: Suppose the green portion of our red-green-blue sequence starts at precision $r_{1}$ and ends at precision $r_{1}+s$, for some $s\geq t$. Let $r_{2}\geq r_{1}+s$ be the minimal precision such that $[r_{2},r]$ is the union of yellow intervals; if no such precision exists, then set $r_{2}=r$.
@@ 5. Reduction to the classical bounds
We note that $e^{\bot}$ and $e$ are Lipschitz-bi-computable, and so $K_{r}(e)=K_{r}(e^{\bot})+O(\log r)$. By a pointwise analog of Marstrand’s projection theorem (which was essentially proven in [19], but is also a result of the projection bound in section 3 arising from a one interval partition), we have
@@ 6. Universal pin sets
In the second subsection, we assume additional regularity in $X$, namely that it is AD regular. Full Ahlfors-David regularity isn’t necessary; it is possible to weaken the assumption slightly. However, we will require AD regularity to simplify the exposition. In the AD regular case, we are guaranteed the existence of points that are extremely regular, with their complexity functions only deviating from $\dim_{H}(x)r$ by a logarithmic term. This enables us to show an optimal effective theorem. We only consider the case that $\dim_{H}(X)>1$ so we can employ a somewhat different reduction, though one that still depends on [24].